Find the exact value of the given functions.
Given tan α = -4/3, α in Quadrant II, and tan β = 24/7, β in Quadrant III, find:
[a] sin(α − β)
[b] cos(α + β)
[c] tan(α − β)

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

This is all covered in great detail on wikipedia - half angle formulas.

Answer by ewatrrr(24785)

tan α = -4/3, α in Quadrant II
sin α = 4/5, cos α = -3/5
...........
tan β = 24/7, β in Quadrant III,
sin β = -24/25, cos β = -7/25
..... Plug and Play
sin(α − β)=
sin(x - y) = sin x cosy - cos x sin y
.......Plug and Play
cos(α + β)=
cos(x - y) = cos x cos y + sin x sin y
.........Plug and Play
tan(x - y) = tan x - tan y)/(1- tan x tan y
tan(α − β) =
..........
(cosx, sinx) Summary Unit Circle
.